If a fair die rolled once comes up 6 you win $100. Otherwise, you lose $15. Would a rational person play this game? Justify your answer, using the definition of E(X).
Answer by stanbon(75887)

If a fair die rolled once comes up 6 you win $100. Otherwise, you lose $15. Would a rational person play this game? Justify your answer, using the definition of E(X).
----
Probability of rolling a 6: 1/6
Probability of not rolling a 6: 5/6
------------------------------------------
E(x) = (1/6)*100 + (5/6)(-15)
---
E(x) = (100 - 75)/6 = 25/6 = $4.17
----
The player can expect to win $4.17 each time he/she plays the game.
===================================
Cheers,
Stan H.